OpenStudy (anonymous):

What is the equation of the line passing through the point (-4, -5) and having a slope of 4?

OpenStudy (anonymous):

y=4x+21

OpenStudy (anonymous):

y=4x-21

OpenStudy (anonymous):

y=4x+11

OpenStudy (anonymous):

y=4x-11

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

y=4x+11 goes up 4

OpenStudy (anonymous):

y=4x+21 goes up 4 too

OpenStudy (igreen):

Plug those into Point-Slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y + 5 = 4(x + 4)\) Distribute 4 into the parenthesis: \(y + 5 = 4x + 16\) Subtract 5 to both sides, what's 16 - 5?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

y=4x+11?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thx